Fermi National Laboratory
 

Accelerator Update for April 20, 2009 - April 22, 2009

 

Monday April 20

The day shift began with Operations monitoring the Tevatron (TeV) store 6988 (110.41E30) and the Recycler stash (288.0E10), with the Antiproton Source stacking (11.8E10), with the Linac, Booster, and Main Injector (MI) operating normally, with the Meson MTest, MiniBooNE, and NuMI experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams.

Operations established beam to MTest at 9:00 AM.

Operations terminated store 6988 at 12:30 PM.

Operations established store 6989 at 2:30 PM with an initial luminosity of 332.62E30.

Operations disabled the MTest beam at 6 PM.

Tuesday April 21

The midnight shift began with Operations monitoring store 6989 (112.53E30) and the stash (281.0E10), with Pbar stacking (18.7E10), and with NuMI and MiniBooNE taking beam.

Operations terminated store 6989 at 4 AM.

Operations established store 6990 at 5:30 AM with an initial luminosity of 329.87E30.

At 2 PM, a Main Injector LCW cooling system tripped off, which inhibited beam to Pbar, NuMI, and MTest through to the end of shift. MiniBooNE users were unaffected by the problem and took protons smoothly all day.

The Main Injector cooling problem was resolved by 4:30 PM. Operations resumed sending beam to Pbar, NuMI, and MTest.

MTest users stopped taking beam at 4:45 PM.

Operations terminated store 6990 at 11:30 PM and put the TeV into HEP shot setup.

Wednesday April 22

The midnight shift began with the TeV in shot setup, with Operations monitoring the stash (384.4E10), with Pbar stacking (42.1E10), and with NuMI and MiniBooNE taking beam.

The HEP shot setup was slowed by a Main Injector damper system problem.

Operations established store 6991 at 3:30 AM with an initial luminosity of 320.92E30.

At 3:45 AM, a Pbar controls crate power supply failed. Operators had the supply replaced by 4:45 AM.

MTest users began taking beam at 7 AM.

The Plans for Wednesday

The plans for today are to stack, stash, and store.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 04/22/2009    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory